40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6598     Entry time: Thu May 3 17:15:38 2012
Author: Koji 
Type: Update 
Category: LSC 
Subject: c1iscaux2 rebooted/burtrestored 

[Jenne/Den/Koji]

We saw some white boxes on the LSC screens.
We found c1iscaux2 is not running.

Once the target was power-cycled, these epics channels are back.
Then c1iscaux2 were burtrestored using the snapshot at 5:07 on 4/16, a day before the power glitch.

ELOG V3.1.3-